OpenStudy (anonymous):

a group of bike riders took a 4 hour trip during the first 3 hours they traveled a total of 50 km during the last hour they traveled only 10 km what was the groups average speed for the entire trip

OpenStudy (anonymous):

S=d/t 50+10=60 60/4=15km/hr